OpenStudy (anonymous):

help please!!!Can someone help me write -2x^2(-5x^2+4x^3) in standard form!!!!! Please

OpenStudy (skullpatrol):

What is the standard form of a polynomial?

OpenStudy (anonymous):

to put the terms with the highest degree first? @skullpatrol

OpenStudy (skullpatrol):

Yes @Klynn30

OpenStudy (skullpatrol):

$$\huge -2x^2(-5x^2+4x^3) $$ recall the distributive property:$$\huge a(b+c)=ab+ac$$

OpenStudy (skullpatrol):

$$\huge a^m\times a^n=a^{m+n}$$To multiply two powers having the same base, you add the exponents.

OpenStudy (skullpatrol):

Does that help @Klynn30 ?